OpenStudy (anonymous):

5(x-4)+9=2x+3(-4+x) x=?

OpenStudy (hoblos):

5x-20 +9 = 2x-12+3x -11=-12 impossible

OpenStudy (anonymous):

5X-11=5X-12

OpenStudy (apoorvk):

no solutions....

OpenStudy (apoorvk):

non-sense basically.. what the above equation is called.